# Combining Spatial and Temporal Abstraction in Planning for Better Generalization

### 17 Jan 2024 v2

@@ 1 Introduction
By making use of imagination and intuition, human conscious planning breaks down long-horizon tasks into more manageable abstract steps, each of which can be narrowed down further. This type of planning attends to important decision points (Sutton et al., 1999) and relevant environmental factors linking the decision points (Tang et al., 2020), thus operating abstractly both in time and in space (Dehaene et al., 2020). In contrast, existing Reinforcement Learning (RL) agents either operate solely based on intuition (model-free methods) or are limited to reasoning over mostly relatively shortsighted plans (model-based methods) Kahneman (2017). The intrinsic limitations constrain the application of RL in real-world under a glass ceiling formed by challenges of longer-term generalization, below the level of human conscious reasoning.
In this paper, we leverage these intuitions to develop a planning agent that automatically decomposes the complex task at hand into smaller subtasks, by constructing abstract “proxy” problems. A proxy problem is represented as a graph where 1) the vertices consist of states proposed by a generative model, corresponding to sparse decision points; and 2) the edges, which define temporally-extended transitions, are constructed by focusing on a small amount of relevant information from the states, using an attention mechanism. Once a proxy problem is constructed and the agent solves it to form a plan, each of the edges defines a new sub-problem, on which the agent will focus next. This divide-and-conquer strategy allows constructing partial solutions that generalize better to new situations, while also giving the agent flexibility to construct abstractions necessary for the problem at hand. Our theoretical analysis establishes guarantees on the quality of the solution to the overall problem.
@@ 3.3 Problem 1: Edge Estimation
First, we discuss how to estimate the edges of the proxy problem, given a set of already generated checkpoints. Inspired by conscious information processing in brains (Dehaene et al., 2020), we introduce a local perceptive field selector, $\sigma$, consisting of an attention bottleneck that (soft-)selects the top-$k$ local segments of the full state (e.g. a feature map by a typical convolutional encoder); all segments of the state compete for the $k$ attention slots, i.e. irrelevant aspects of state are discarded, to form a partial state representation (Mott et al., 2019; Tang et al., 2020; Zhao et al., 2021; Alver & Precup, 2023). We provide an example in Fig. 2 (see purple parts). Through $\sigma$, the auxiliary estimators, to be discussed soon, force the bottleneck mechanism to promote aspects relevant to the local estimation of connections between the checkpoints. The rewards and discounts are then estimated from the partial state $\sigma(S)$, based on the agent’s behavior.
@@ 4 Related Works & Discussions
Temporal Abstraction. Constraining options, Skipper learns the options targeting certain “outcomes”, which dodges the difficulties of option collapse (Bacon et al., 2017) and option outcome modelling by design. The constraints indeed shift the difficulties to generator learning (Silver & Ciosek, 2012; Tang & Salakhutdinov, 2019). We expect this to entail benefits where states are easy to learn and generate, and / or in stochastic environments where the outcomes of unconstrained options are difficult to learn. Constraining options has also been investigated in Sharma et al. (2019) via a predictability-oriented intrinsic reward in an unsupervised setting.
@@
Spatial Abstraction is different from “state abstraction” (Sacerdoti, 1974; Knoblock, 1994), which evolved to be a more general concept that embraces mainly the aspect of state aggregation, i.e. state space partitioning (Li et al., 2006). Defined to be more precise in what we mean, spatial abstraction focuses on the within-state partial selection of environment information for decision-making, an idea traceable to early hand-coded proof-of-concepts concepts proposed in e.g. Dietterich (2000). It corresponds naturally to the intuition that state representations should contain useful aspects of the environment, while not all aspects are useful for a particular intent. The approaches of spatio-temporal abstractions can be seen to a degree orthogonal to the traditional approaches of state-action abstractions (Shah et al., 2021). Until only recently, attention mechanisms had primarily been used to construct state representations in model-free agents for sample efficiency purposes, without the focus on generalization (Mott et al., 2019; Manchin et al., 2019; Tang et al., 2020). In Fu et al. (2021); Zadaianchuk et al. (2020); Shah et al. (2021), $3$ more recent model-based approaches, spatial abstractions are attempted to remove visual distractors. Concurrently, emphasizing on generalization, Zhao et al. (2021) used spatially-abstract partial states in decision-time planning. The authors proposed an attention bottleneck to dynamically select a subset of environmental entities during the atomic-step forward simulation, without explicit goals provided as in Zadaianchuk et al. (2020). Skipper’s checkpoint transition is a step-up approach, where we show that spatial abstraction, an overlooked missing flavor, is as crucial for longer-term planning as temporal abstraction (Konidaris & Barto, 2009). Shah et al. (2021) attempted longer-horizon reasoning with an empirical framework that leverages state-action abstractions, yet with pre-specified set of skills rather than learned like in this work, which provides a theory-regulated framework.
@@
Task Abstraction via Goal Composition In the early work McGovern & Barto (2001), the authors suggested to use bottlenecks of state spaces as subgoals to abstract given tasks into manageable steps. Nair et al. (2018) uses generative models to imagine subgoals while Eysenbach et al. (2019) search directly on the experience replay. In Kim et al. (2021), promising states to explore are generated and selected with shortest-path algorithms. Similar ideas have been attempted for guided exploration (Erraqabi et al., 2021; Kulkarni et al., 2016). Similar to Hafner et al. (2022), Czechowski et al. (2021) generate $k$-th step ahead subgoals for complex reasoning tasks, while Bagaria et al. (2021) augments the search graph by states reached $k$-th step ahead. Nasiriany et al. (2019); Xie et al. (2020); Shi et al. (2022) employ CEM to plan a chain of subgoals towards the task goal (Rubinstein, 1997b). Skipper utilizes proxy problems, which achieve the abstraction of the given tasks via spatio-temporal abstractions (Bagaria et al., 2021). Checkpoints can be seen as sub-goals that generalize the notion of “landmarks” or “waypoints” in Sutton et al. (1999); Dietterich (2000); Savinov et al. (2018). Zhang et al. (2021) explored a framework of using latent landmark graphs as high-level guidance, where the landmarks are selected via a sparsification procedure that uses a weighted sum in the latent space to compose subgoals. In comparison, our checkpoint pruning selects a subset of generated states, which is less prone to issues created by weighted sums.
@@ 5 Experiments
As introduced in Sec. 2, our first goal is to test the zero-shot generalization ability of trained agents. To fully understand the results, it is necessary to have precise control of the difficulty of the training and evaluation tasks. Also, to validate if the empirical performance of our agents matches the formal analyses (Thm. 1), we need to know how close to the (optimal) ground truth our edge estimations and checkpoint policies are. These goals lead to the need for environments whose ground truth information (optimal policies, true distances between checkpoints, etc) can be computed. Thus, we base our experimental setting on the MiniGrid-BabyAI framework Chevalier-Boisvert et al. (2018b; a); Hui et al. (2020). Specifically, we build on the experiments used in Zhao et al. (2021); Alver & Precup (2022): the agent needs to navigate to the goal from its initial state in gridworlds filled with terminal lava traps generated randomly according to a difficulty parameter, which controls their density. During evaluation, the agent is always spawned at the opposite side from the goals. During training, the agent’s position is uniformly initialized to speed up training. We provide results for non-uniform training initialization in the Appendix.
@@ 6 Conclusion & Future Work
We proposed, analyzed and validated our HP framework, Skipper, which generalizes its learned skills better than the compared methods, due to its spatio-temporal abstractions. In this work, we generated checkpoints at random by sampling the partial description space. Despite the pruning mechanisms, the generated checkpoints, and thus temporal abstraction, do not prioritize the predictable, important states that matter to form a meaningful plan (Şimşek & Barto, 2004). We would like to continue investigating the possibilities along this line. Additionally, we would like to explore other environments where the accuracy assumption (in Thm. 1) can meaningfully hold, i.e. beyond sparse reward cases.
@@ E.7 Hyperparameter Tuning
Model-free Baseline Architecture The baseline architecture (distributional, Double DQN) was heavily influenced by the architecture used in Zhao et al. (2021), which demonstrated success on similar but smaller-scale experiments ($8\times 8$). The difference is that while they used computationally heavy components such as transformer layers on a set-based representation, we replaced them with a simpler and effective local perception component. We validated our model-free baseline performance on the tasks proposed in Zhao et al. (2021).

### 04 Feb 2024 v3

# Consciousness-Inspired Spatio-Temporal Abstraction for Better Generalization

@@ 1 Introduction
By attending to relevant aspects in both space and time, human conscious planning breaks down long-horizon tasks into more manageable abstract steps, each of which can be narrowed down further. Stemming from consciousness in the first sense (C1) (Dehaene et al., 2020), this type of planning focuses attention on mostly the important decision points (Sutton et al., 1999) and relevant environmental factors linking the decision points (Tang et al., 2020), thus operating abstractly both in time and in space. In contrast, existing Reinforcement Learning (RL) agents either operate solely based on intuition (model-free methods) or are limited to reasoning over mostly relatively shortsighted plans (model-based methods) Kahneman (2017). The intrinsic limitations constrain the application of RL in real-world under a glass ceiling formed by challenges of longer-term generalization, below the level of human conscious reasoning.
Building on our previous work on conscious planning (Zhao et al., 2021), we take inspirations to develop a planning agent that automatically decomposes the complex task at hand into smaller subtasks, by constructing abstract “proxy” problems. A proxy problem is represented as a graph where 1) the vertices consist of states proposed by a generative model, corresponding to sparse decision points; and 2) the edges, which define temporally-extended transitions, are constructed by focusing on a small amount of relevant information from the states, using an attention mechanism. Once a proxy problem is constructed and the agent solves it to form a plan, each of the edges defines a new sub-problem, on which the agent will focus next. This divide-and-conquer strategy allows constructing partial solutions that generalize better to new situations, while also giving the agent flexibility to construct abstractions necessary for the problem at hand. Our theoretical analysis establishes guarantees on the quality of the solution to the overall problem.
@@ 3.3 Problem 1: Edge Estimation
First, we discuss how to estimate the edges of the proxy problem, given a set of already generated checkpoints. Inspired by conscious information processing in brains (Dehaene et al., 2020) and existing approach in Sylvain et al. (2019), we introduce a local perceptive field selector, $\sigma$, consisting of an attention bottleneck that (soft-)selects the top-$k$ local segments of the full state (e.g. a feature map by a typical convolutional encoder); all segments of the state compete for the $k$ attention slots, i.e. irrelevant aspects of state are discarded, to form a partial state representation (Mott et al., 2019; Tang et al., 2020; Zhao et al., 2021; Alver & Precup, 2023). We provide an example in Fig. 2 (see purple parts). Through $\sigma$, the auxiliary estimators, to be discussed soon, force the bottleneck mechanism to promote aspects relevant to the local estimation of connections between the checkpoints. The rewards and discounts are then estimated from the partial state $\sigma(S)$, based on the agent’s behavior.
@@ 4 Related Works & Discussions
Temporal Abstraction. Resembling attention in human consciousness, choosing a checkpoint target is a selection towards certain decision points in the dimension of time, i.e. a form of temporal abstraction. Constraining options, Skipper learns the options targeting certain “outcomes”, which dodges the difficulties of option collapse (Bacon et al., 2017) and option outcome modelling by design. The constraints indeed shift the difficulties to generator learning (Silver & Ciosek, 2012; Tang & Salakhutdinov, 2019). We expect this to entail benefits where states are easy to learn and generate, and / or in stochastic environments where the outcomes of unconstrained options are difficult to learn. Constraining options was also investigated in Sharma et al. (2019) in an unsupervised setting.
@@
Spatial Abstraction is different from “state abstraction” (Sacerdoti, 1974; Knoblock, 1994), which evolved to be a more general concept that embraces mainly the aspect of state aggregation, i.e. state space partitioning (Li et al., 2006). Spatial abstraction, defined to capture the behavior of attention in conscious planning in the spatial dimensions, focuses on the within-state partial selection of the environmental state for decision-making. It corresponds naturally to the intuition that state representations should contain useful aspects of the environment, while not all aspects are useful for a particular intent. Efforts toward spatial abstraction are traceable to early hand-coded proof-of-concepts proposed in e.g. Dietterich (2000). Until only recently, attention mechanisms had primarily been used to construct state representations in model-free agents for sample efficiency purposes, without the focus on generalization (Mott et al., 2019; Manchin et al., 2019; Tang et al., 2020). In Fu et al. (2021); Zadaianchuk et al. (2020); Shah et al. (2021), $3$ more recent model-based approaches, spatial abstractions are attempted to remove visual distractors. Concurrently, emphasizing on generalization, our previous work (Zhao et al., 2021) used spatially-abstract partial states in decision-time planning. We proposed an attention bottleneck to dynamically select a subset of environmental entities during the atomic-step forward simulation, without explicit goals provided as in Zadaianchuk et al. (2020). Skipper’s checkpoint transition is a step-up from our old approach, where we now show that spatial abstraction, an overlooked missing flavor, is as crucial for longer-term planning as temporal abstraction (Konidaris & Barto, 2009).
@@
Task Abstraction via Goal Composition The early work McGovern & Barto (2001) suggested to use bottleneck states as subgoals to abstract given tasks into manageable steps. Nair et al. (2018); Florensa et al. (2018) use generative model to imagine subgoals while Eysenbach et al. (2019) search directly on the experience replay. In Kim et al. (2021), promising states to explore are generated and selected with shortest-path algorithms. Similar ideas have been attempted for guided exploration (Erraqabi et al., 2021; Kulkarni et al., 2016). Similar to Hafner et al. (2022), Czechowski et al. (2021) generate fixed-steps ahead subgoals for reasoning tasks, while Bagaria et al. (2021) augments the search graph by states reached fixed-steps ahead. Nasiriany et al. (2019); Xie et al. (2020); Shi et al. (2022) employ CEM to plan a chain of subgoals towards the task goal (Rubinstein, 1997b). Skipper utilizes proxy problems to abstract the given tasks via spatio-temporal abstractions (Bagaria et al., 2021). Checkpoints can be seen as sub-goals that generalize the notion of “landmarks” or “waypoints” in Sutton et al. (1999); Dietterich (2000); Savinov et al. (2018). Zhang et al. (2021) used latent landmark graphs as high-level guidance, where the landmarks are sparsified with weighted sums in the latent space to compose subgoals. In comparison, our checkpoint pruning selects a subset of generated states, which is less prone to issues created by weighted sums.
@@ 5 Experiments
As introduced in Sec. 2, our first goal is to test the zero-shot generalization ability of trained agents. To fully understand the results, it is necessary to have precise control of the difficulty of the training and evaluation tasks. Also, to validate if the empirical performance of our agents matches the formal analyses (Thm. 1), we need to know how close to the (optimal) ground truth our edge estimations and checkpoint policies are. These goals lead to the need for environments whose ground truth information (optimal policies, true distances between checkpoints, etc) can be computed. Thus, we base our experimental setting on the MiniGrid-BabyAI framework Chevalier-Boisvert et al. (2018b; a); Hui et al. (2020). Specifically, we build on the experiments used in our previous works Zhao et al. (2021); Alver & Precup (2022): the agent needs to navigate to the goal from its initial state in gridworlds filled with terminal lava traps generated randomly according to a difficulty parameter, which controls their density. During evaluation, the agent is always spawned at the opposite side from the goals. During training, the agent’s position is uniformly initialized to speed up training. We provide results for non-uniform training initialization in the Appendix.
@@ 6 Conclusion & Future Work
Building on previous work on spatial abstraction (Zhao et al., 2021), we proposed, analyzed and validated Skipper, which generalizes its learned skills better than the compared methods, due to its combinned spatio-temporal abstractions. A major unsatisfactory aspect of this work is that we generated checkpoints at random by sampling the partial description space. Despite the pruning mechanisms, the generated checkpoints, and thus temporal abstraction, do not prioritize the predictable, important states that matter to form a meaningful plan (Şimşek & Barto, 2004). We would like to continue investigating the possibilities along this line. Additionally, we would like to explore other environments where the accuracy assumption (in Thm. 1) can meaningfully hold, i.e. beyond sparse reward cases.
@@ E.7 Hyperparameter Tuning
Model-free Baseline Architecture The baseline architecture (distributional, Double DQN) was heavily influenced by the architecture used in the previous work (Zhao et al., 2021), which demonstrated success on similar but smaller-scale experiments ($8\times 8$). The difference is that while then we used computationally heavy components such as transformer layers on a set-based representation, we replaced them with a simpler and effective local perception component. We validated our model-free baseline performance on the tasks proposed in Zhao et al. (2021).